# China Regional 3km Downscaling Based on Residual Corrective Diffusion Model

### v3

@@ 1 Introduction
Gridded meteorological forecasts are important in various fields such as transportation, energy sector, agriculture, and scientific research. In particular, high spatial resolution forecasts are crucial for local studies and risk assessment. Traditionally, global gridded forecasts are obtained from numerical weather prediction models. Generating km-resolution forecasts is challenging for numerical weather prediction models due to the limitation of computation time. Currently, there are also data-driven deep learning models that generate global forecasts [bi2022pangu, lam2023learning, pathak2022fourcastnet, chen2023fuxi]. However, up to now, there are few global high-resolution gridded reanalysis data that can be used to train such data-driven models. Instead of getting global high-resolution forecasts, a practical alternative is to get regional high-resolution forecasts from the low-resolution output of a global model by downscaling.
@@
Many machine learning models have been applied for statistical downscaling, such as multiple linear regression [schoof2001downscaling], support vector machine [chen2010downscaling], random forest [davy2010statistical], and artificial neural networks [laddimath2019artificial]. Among these models, artificial neural networks, which have evolved into deep learning [sun2024deep], are likely the most promising approach.
@@
The fast development in deep learning over the past decade has led to numerous active research areas, including super-resolution in computer vision. Various high-performance super-resolution models have been proposed, such as Generative Adversarial Networks (GANs) and diffusion models. Super-resolution is similarity to downscaling: the input for both is a low-resolution grid, and the output is a high-resolution grid. However, there is also a difference between super-resolution and downscaling: the goal of super-resolution is to generate visually realistic images, while meteorological downscaling must ensure accuracy and physical consistency. Given the similarity between super-resolution and downscaling, many researchers have investigated the application of such super-resolution models to downscaling [watt2024generative, addison2022machine]. In parallel, there are also works that developed specific neural network structures for downscaling [wu2024gsdnet].
@@
This work investigates a diffusion-based downscaling model named Corrective Diffusion (CorrDiff) [mardani2025residual]. CorrDiff is a two-step approach that includes the training of a regression model and the training of a diffusion model to improve the predictions of the regression model. In [mardani2025residual], CorrDiff is applied to the Taiwan region, the resolution of the inputs is 25km and the resolution of the outputs is 2km. The size of the 2km high-resolution grid is 448 $\times$ 448. In this study, we apply CorrDiff on the China region, the resolutions of the inputs and the outputs are 25km and 3km respectively. The size of our 3km high-resolution grid is 1600 $\times$ 2400, which is nearly 20 times the size in [mardani2025residual]. Our models are trained on reanalysis data, including 25km ECMWF Reanalysis v5 (ERA5) [hersbach2020era5] (as low-resolution inputs of the downscaling models) and 3km reanalysis data (as high-resolution labels) that are produced by China Meteorological Administration Regional Reanalysis Atmospheric System (CMA-RRA). In contrast to [mardani2025residual], which focuses mainly on surface variables, in this work multiple variables are considered for downscaling, including surface variables and variables at six pressure levels. The prediction of radar composite reflectivity is also investigated. Different combinations of input and output variables are examined in order to understand the intervariable dependencies in the downscaling task. By connecting our downscaling models to global forecasts, 3km regional forecasts for the China region are obtained, which are evaluated through comparison with CMA-MESO. CMA-MESO is a high-resolution regional numerical weather prediction model of the China Meteorological Administration (CMA) that generates 3km and 1km resolution forecasts of the China region. Two global forecasts are considered: CMA-GFS, which is an operational global model of CMA, and SFF, which is a deep learning-based weather model. The experimental results demonstrate that, for the mean absolute error (MAE), our forecasts generally outperform those of CMA-MESO for the target variables. Our assessment of the uncertainty estimated by CorrDiff shows that, for any downscaled variable, the uncertainty is correlated with the accuracy. For the prediction of radar reflectivity, our results show that deterministic models usually have a severe over-smoothing problem, while CorrDiff can generate realistic small-scale features that are similar to reanalysis data.
@@ 2 Problem Formulation
In this work, we also use diffusion models. Generally, a diffusion model involves two processes: noising and denoising. Noising is the process that progressively adds noise to a clean image until the image becomes random noise. Denoising is the process that transforms random noise into a structured image. Basically, the training of a diffusion model can be considered as the training of a denoising model. In the case of using diffusion models for downscaling, the low-resolution data are also fed into the diffusion models. Such models are also known as conditional diffusion models. The overall inference process of a conditional diffusion model for downscaling can be represented by a function $f:\mathbb{R}^{c_{in}\times p\times q},E\mapsto\mathbb{R}^{c_{out}\times m\times n}$, which is a conditional sampling function of the probability of the 3km resolution data given a 25km resolution input. $E$ represents the set of random noise that follows a certain distribution (which normally follows the normal distribution). Note that this function $f$ is highly complex because the denoising process contains multiple steps that progressively decrease the noise level. For more details on diffusion models, see [ho2020denoising].
@@ 3 Corrective Diffusion Model
This section briefly introduces the Corrective Diffusion Model (CorrDiff), for more details see [mardani2025residual]. The training of CorrDiff has two steps. In the first step, we train a regression model (UNet), denoted as $f$, to minimize a loss function between its predictions $f(bilinear(x_{i}))$ and the targets $y_{i}$ for all $i\in\{1,...,N\}$. In the second step, we train an Elucidated Diffusion Model (EDM) [karras2022elucidating] to correct the predictions made by $f$. We denote the overall inference process of EDM by a function $g$. After training, the CorrDiff prediction $\hat{y}_{i}$ for the target $y_{i}$ is given by $\hat{y}_{i}=g(bilinear(x_{i}),f(bilinear(x_{i})),\epsilon)+f(bilinear(x_{i}))$, where $\epsilon$ is a random noise sampled from a standard normal distribution. By sampling multiple random noises, a CorrDiff model generates an ensemble of predictions, thereby producing a probabilistic prediction of $y_{i}$.
Following [mardani2025residual], we use a specific architecture of UNet [karras2022elucidating] for both the regression (first step) and diffusion (second step) networks. This architecture has 6 encoder and decoder layers, and it incorporates attention mechanisms and residual connections within its structure.
Fig. 1 shows the overall architecture of one of our trained CorrDiff models. Our regression networks differ slightly from the models in [mardani2025residual]: for the downscaling of a variable, assuming that $x_{i}^{u}$ and $y_{i}^{v}$ correspond to this variable (where $x_{i}^{u}\in\mathbb{R}^{p\times q}$ is a channel of $x_{i}$, and $y_{i}^{v}\in\mathbb{R}^{m\times n}$ is a channel of $y_{i}$), instead of directly predicting $y_{i}^{v}$, we predict the residual $y_{i}^{v}-bilinear(x_{i}^{u})$, because our experimental results show that this residual learning strategy could accelerate convergence speed and improve accuracy for the downscaling task. In contrast to [mardani2025residual] which only considers the downscaling of surface variables, this work also investigates the downscaling of pressure level variables. An additional input (3km resolution orography) is included in our models as well. The implementation of our models is based on the code provided in [mardani2025residual].
@@
As stated in [mardani2025residual], the development of CorrDiff was motivated by the limitations of using conditional diffusion models for downscaling. It is assumed that there is a significant distribution shift between low-resolution and high-resolution data that hinders the learning process. To avoid this problem, the generation is decomposed into two steps. The first step aims to predict the conditional mean using regression (UNet), and the second step learns a correction using a diffusion model.
@@ 4.1 Baseline Model: CMA-MESO
The China Meteorological Administration Mesoscale (CMA-MESO) model is a high-resolution regional model developed by the China Meteorological Administration Earth System Modeling and Prediction Centre (CEMC)[liping2022key]. The model is based on a fully compressible non-hydrostatic dynamical core that uses a semi-implicit semi-Lagrangian discretization scheme. It covers the region of 10–60.1°N, 70–145°E. Currently, CMA-MESO has two horizontal resolutions: $\times 0.01$°($\approx$ 1km) and $\times 0.03$°($\approx$ 3km). This study only considers the 3km version, as the required reanalysis data are currently available at this resolution. We select CMA-MESO as the baseline model because the long-term objective of this work is to develop a deep learning-based alternative to CMA-MESO for specific tasks.
@@ 5.1.2 CorrDiff Models
The CorrDiff models that correspond to Regression 1, 2, and 4 are denoted as CorrDiff 1, 2, and 4, respectively. Their MAE scores on the validation set are presented in Table 2, where Regression 1, 2, and 4 were trained for 91, 120 and 35 epochs, respectively, and CorrDiff 1, 2, and 4 were trained for 285, 100 and 59 epochs, respectively. All reported MAE scores for the CorrDiff models are computed using a single sample. Our results indicate that the MAE scores of the regression models are consistently lower than those of the CorrDiff models. In fact, the degradation in MAE of the CorrDiff models compared to that of the regression models is also reported in [mardani2025residual]. As stated in [mardani2025residual], the degradation is expected as the diffusion models optimize the Kullback-Leibler divergence as opposed to the regression models that minimize the MSE loss. Following [mardani2025residual], we also compute the continuous ranked probability score (CRPS) [hersbach2000decomposition], see Table 3. CRPS can be considered as a generalization of the mean absolute error for probability prediction. Since the computation of CRPS is slow with our current implementation (primarily due to the large grid size), we select a more restricted set of time points for its calculation: 00:00 and 12:00 UTC on the first day of each month in 2023 and we only focus on CorrDiff 4. As expected, the CRPS values are generally the lowest.
@@ 5.2.1 SFF
Sphere Fusion Forecast (SFF) is a data-driven deep learning-based weather model developed from Spherical Fourier Neural Operators (SFNO) [bonev2023spherical]. Compared to SFNO, two major improvements are made in SFF: the up-sampling and down-sampling operators between the SFNO blocks are added, allowing the initial and final stages of the SFNO block chain to handle broader frequency spectra, while the middle layers focus on relatively low-frequency information; a Vision Transformer–like architecture between the encoder and decoder is introduced as the skip connection, which improves the model’s capacity for local feature learning, producing more robust and accurate forecasts.
@@ 5.2.2 CMA-GFS
China Meteorological Administration-Global Forecast System (CMA-GFS) is a global model developed by the CEMC. It comprises a semi-implicit semi-Lagrangian (SISL) non-hydrostatic dynamical core, a physical parameterization package, and a four-dimensional variational data assimilation system [chen2008new]. Currently, the horizontal resolution of CMA-GFS is $\times 0.125$°($\approx$ 12.5km).
@@ 6 Conclusion
  • A challenge remains in the interpretability of deep learning models [zhang2021survey]. Various methods (for example, gradient-based approaches [simonyan2013deep, sundararajan2017axiomatic]) have been developed for computer vision tasks such as image classification. These techniques could be adapted to elucidate the predictions of CorrDiff by incorporating physical principles.

### v4

@@ 1 Introduction
Gridded meteorological forecasts are important in various fields such as transportation, energy sector, agriculture, and scientific research. In particular, high spatial resolution forecasts are crucial for local studies and risk assessment. Traditionally, global gridded forecasts are obtained from numerical weather prediction models. Generating km-resolution forecasts is challenging for numerical weather prediction models due to the limitation of computation time. Currently, there are also data-driven deep learning models that generate global forecasts [2, 13, 16, 6]. However, up to now, there are few global high-resolution gridded reanalysis data that can be used to train such data-driven models. Instead of getting global high-resolution forecasts, a practical alternative is to get regional high-resolution forecasts from the low-resolution output of a global model by downscaling.
@@
Many machine learning models have been applied for statistical downscaling, such as multiple linear regression [17], support vector machine [5], random forest [7], and artificial neural networks [12]. Among these models, artificial neural networks, which have evolved into deep learning [19], are likely the most promising approach.
@@
The fast development in deep learning over the past decade has led to numerous active research areas, including super-resolution in computer vision. Various high-performance super-resolution models have been proposed, such as Generative Adversarial Networks (GANs) and diffusion models. Super-resolution is similarity to downscaling: the input for both is a low-resolution grid, and the output is a high-resolution grid. However, there is also a difference between super-resolution and downscaling: the goal of super-resolution is to generate visually realistic images, while meteorological downscaling must ensure accuracy and physical consistency. Given the similarity between super-resolution and downscaling, many researchers have investigated the application of such super-resolution models to downscaling [21, 1]. In parallel, there are also works that developed specific neural network structures for downscaling [22].
@@
This work investigates a diffusion-based downscaling model named Corrective Diffusion (CorrDiff) [15]. CorrDiff is a two-step approach that includes the training of a regression model and the training of a diffusion model to improve the predictions of the regression model. In [15], CorrDiff is applied to the Taiwan region, the resolution of the inputs is 25km and the resolution of the outputs is 2km. The size of the 2km high-resolution grid is 448 $\times$ 448. In this study, we apply CorrDiff on the China region, the resolutions of the inputs and the outputs are 25km and 3km respectively. The size of our 3km high-resolution grid is 1600 $\times$ 2400, which is nearly 20 times the size in [15]. Our models are trained on reanalysis data, including 25km ECMWF Reanalysis v5 (ERA5) [8] (as low-resolution inputs of the downscaling models) and 3km reanalysis data (as high-resolution labels) that are produced by China Meteorological Administration Regional Reanalysis Atmospheric System (CMA-RRA). In contrast to [15], which focuses mainly on surface variables, in this work multiple variables are considered for downscaling, including surface variables and variables at six pressure levels. The prediction of radar composite reflectivity is also investigated. Different combinations of input and output variables are examined in order to understand the intervariable dependencies in the downscaling task. By connecting our downscaling models to global forecasts, 3km regional forecasts for the China region are obtained, which are evaluated through comparison with CMA-MESO. CMA-MESO is a high-resolution regional numerical weather prediction model of the China Meteorological Administration (CMA) that generates 3km and 1km resolution forecasts of the China region. Two global forecasts are considered: CMA-GFS, which is an operational global model of CMA, and SFF, which is a deep learning-based weather model. The experimental results demonstrate that, for the mean absolute error (MAE), our forecasts generally outperform those of CMA-MESO for the target variables. Our assessment of the uncertainty estimated by CorrDiff shows that, for any downscaled variable, the uncertainty is correlated with the accuracy. For the prediction of radar reflectivity, our results show that deterministic models usually have a severe over-smoothing problem, while CorrDiff can generate realistic small-scale features that are similar to reanalysis data.
@@ 2 Problem Formulation
In this work, we also use diffusion models. Generally, a diffusion model involves two processes: noising and denoising. Noising is the process that progressively adds noise to a clean image until the image becomes random noise. Denoising is the process that transforms random noise into a structured image. Basically, the training of a diffusion model can be considered as the training of a denoising model. In the case of using diffusion models for downscaling, the low-resolution data are also fed into the diffusion models. Such models are also known as conditional diffusion models. The overall inference process of a conditional diffusion model for downscaling can be represented by a function $f:\mathbb{R}^{c_{in}\times p\times q},E\mapsto\mathbb{R}^{c_{out}\times m\times n}$, which is a conditional sampling function of the probability of the 3km resolution data given a 25km resolution input. $E$ represents the set of random noise that follows a certain distribution (which normally follows the normal distribution). Note that this function $f$ is highly complex because the denoising process contains multiple steps that progressively decrease the noise level. For more details on diffusion models, see [10].
@@ 3 Corrective Diffusion Model
This section briefly introduces the Corrective Diffusion Model (CorrDiff), for more details see [15]. The code implementation is based on PhysicsNeMo (https://github.com/NVIDIA/physicsnemo). The training of CorrDiff has two steps. In the first step, we train a regression model (UNet), denoted as $f$, to minimize a loss function between its predictions $f(bilinear(x_{i}))$ and the targets $y_{i}$ for all $i\in\{1,...,N\}$. In the second step, we train an Elucidated Diffusion Model (EDM) [11] to correct the predictions made by $f$. We denote the overall inference process of EDM by a function $g$. After training, the CorrDiff prediction $\hat{y}_{i}$ for the target $y_{i}$ is given by $\hat{y}_{i}=g(bilinear(x_{i}),f(bilinear(x_{i})),\epsilon)+f(bilinear(x_{i}))$, where $\epsilon$ is a random noise sampled from a standard normal distribution. By sampling multiple random noises, a CorrDiff model generates an ensemble of predictions, thereby producing a probabilistic prediction of $y_{i}$.
Following [15], we use a specific architecture of UNet [11] for both the regression (first step) and diffusion (second step) networks. This architecture has 6 encoder and decoder layers, and it incorporates attention mechanisms and residual connections within its structure.
Fig. 1 shows the overall architecture of one of our trained CorrDiff models. Our regression networks differ slightly from the models in [15]: for the downscaling of a variable, assuming that $x_{i}^{u}$ and $y_{i}^{v}$ correspond to this variable (where $x_{i}^{u}\in\mathbb{R}^{p\times q}$ is a channel of $x_{i}$, and $y_{i}^{v}\in\mathbb{R}^{m\times n}$ is a channel of $y_{i}$), instead of directly predicting $y_{i}^{v}$, we predict the residual $y_{i}^{v}-bilinear(x_{i}^{u})$, because our experimental results show that this residual learning strategy could accelerate convergence speed and improve accuracy for the downscaling task. In contrast to [15] which only considers the downscaling of surface variables, this work also investigates the downscaling of pressure level variables. An additional input (3km resolution orography) is included in our models as well. The implementation of our models is based on the code provided in [15].
@@
As stated in [15], the development of CorrDiff was motivated by the limitations of using conditional diffusion models for downscaling. It is assumed that there is a significant distribution shift between low-resolution and high-resolution data that hinders the learning process. To avoid this problem, the generation is decomposed into two steps. The first step aims to predict the conditional mean using regression (UNet), and the second step learns a correction using a diffusion model.
@@ 4.1 Baseline Model: CMA-MESO
The China Meteorological Administration Mesoscale (CMA-MESO) model is a high-resolution regional model developed by the China Meteorological Administration Earth System Modeling and Prediction Centre (CEMC)[14]. The model is based on a fully compressible non-hydrostatic dynamical core that uses a semi-implicit semi-Lagrangian discretization scheme. It covers the region of 10–60.1°N, 70–145°E. Currently, CMA-MESO has two horizontal resolutions: $\times 0.01$°($\approx$ 1km) and $\times 0.03$°($\approx$ 3km). This study only considers the 3km version, as the required reanalysis data are currently available at this resolution. We select CMA-MESO as the baseline model because the long-term objective of this work is to develop a deep learning-based alternative to CMA-MESO for specific tasks.
@@ 5.1.2 CorrDiff Models
The CorrDiff models that correspond to Regression 1, 2, and 4 are denoted as CorrDiff 1, 2, and 4, respectively. Their MAE scores on the validation set are presented in Table 2, where Regression 1, 2, and 4 were trained for 91, 120 and 35 epochs, respectively, and CorrDiff 1, 2, and 4 were trained for 285, 100 and 59 epochs, respectively. All reported MAE scores for the CorrDiff models are computed using a single sample. Our results indicate that the MAE scores of the regression models are consistently lower than those of the CorrDiff models. In fact, the degradation in MAE of the CorrDiff models compared to that of the regression models is also reported in [15]. As stated in [15], the degradation is expected as the diffusion models optimize the Kullback-Leibler divergence as opposed to the regression models that minimize the MSE loss. Following [15], we also compute the continuous ranked probability score (CRPS) [9], see Table 3. CRPS can be considered as a generalization of the mean absolute error for probability prediction. Since the computation of CRPS is slow with our current implementation (primarily due to the large grid size), we select a more restricted set of time points for its calculation: 00:00 and 12:00 UTC on the first day of each month in 2023 and we only focus on CorrDiff 4. As expected, the CRPS values are generally the lowest.
@@ 5.2.1 SFF
Sphere Fusion Forecast (SFF) is a data-driven deep learning-based weather model developed from Spherical Fourier Neural Operators (SFNO) [3]. Compared to SFNO, two major improvements are made in SFF: the up-sampling and down-sampling operators between the SFNO blocks are added, allowing the initial and final stages of the SFNO block chain to handle broader frequency spectra, while the middle layers focus on relatively low-frequency information; a Vision Transformer–like architecture between the encoder and decoder is introduced as the skip connection, which improves the model’s capacity for local feature learning, producing more robust and accurate forecasts.
@@ 5.2.2 CMA-GFS
China Meteorological Administration-Global Forecast System (CMA-GFS) is a global model developed by the CEMC. It comprises a semi-implicit semi-Lagrangian (SISL) non-hydrostatic dynamical core, a physical parameterization package, and a four-dimensional variational data assimilation system [4]. Currently, the horizontal resolution of CMA-GFS is $\times 0.125$°($\approx$ 12.5km).
@@ 6 Conclusion
  • A challenge remains in the interpretability of deep learning models [23]. Various methods (for example, gradient-based approaches [18, 20]) have been developed for computer vision tasks such as image classification. These techniques could be adapted to elucidate the predictions of CorrDiff by incorporating physical principles.